OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (the_fizicx99):

Apply distributive property, 3(x+1) = -2(x-1) - 4 3x + 3 = -2x + 2 - 4 3x + 3 = -2x -2 -3 -3 3x = -2x -5 +2x +2x 5x = -5 /5 /5 x = -1

OpenStudy (anonymous):

@tHe_FiZiCx99 Thanks

OpenStudy (anonymous):

good job @tHe_FiZiCx99

OpenStudy (the_fizicx99):

thnx